Grade 12Algebra

Give first the step you will use to separate the variable and then solve the equation:
y + 4 = – 4

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

We have to subtract 4 to both the side of given equation,
Then we get,
=y +4 - 4 = - 4 - 4
= y = - 8

Thanks